# Robustness-Aware 3D Object Detection in Autonomous Driving: A Review and Outlook

### v1

@@ IV-B2 Model Structure
Most voxel-based detectors consist of two fundamental components: voxel-based data representation and voxel-based neural networks. Specifically, there are primarily three major types of neural networks of voxel-based methods: 1) 2D CNNs used for processing BEV feature maps and pillars. 2) 3D CNNs for processing voxels. 3) Transformers for handling both voxels and pillars.
2D CNN
: 2D CNN is primarily used to detect 3D objects from a bird’s-eye view perspective, including processing BEV (Bird’s Eye View) feature maps and pillars [118, 183, 119, 181, 135]. Specifically, the 2D CNN used for processing BEV feature maps often come from well-developed 2D object detection networks, such as Darknet [415], ResNet [416], FPN [402], and RPN [401]. Some early voxel-based works drew inspiration from mature ideas in 2D detection, e.g. Voxel-FPN [115]. One significant advantage of 2D CNN compared to 3D CNN is its faster speed. However, due to its difficulty in capturing spatial relationships and shape information, 2D CNN typically exhibits lower accuracy.
3D Sparse CNN
: 3D Sparse CNN consists of two core operators: sparse convolution and submanifold convolution [417], which ensure that the convolutional operation is performed only on non-empty voxels. SECOND [109] implements efficient computation of sparse convolution [417] and submanifold convolution [418] operators to gain fast inference speed by constructing a hash table. Due to its outstanding performance, it’s followed by [112, 125, 113, 250, 104]. However, the limited receptive field of 3D Sparse CNNs, leading to information truncation, restricts the model’s feature extraction capabilities. Meanwhile, the sparse representation of features makes it challenging for the model to capture fine-grained object boundaries and detailed information. To optimize these issues, main optimization strategies have emerged: 1) Expanding the model’s receptive field. Some methods [130, 129] extend the concept of large kernel convolution from 2D to 3D space or introduce additional downsampling layers in the model [112]. 2) Combining sparse and dense representations. Methods in this category typically utilize dense prediction heads to prevent information loss [125, 113, 108, 109, 372] or retrieve lost 3D information from the detection process [113, 168, 372, 180, 102, 125], or they add additional auxiliary tasks to the model [372, 201, 419, 103, 104, 151]. Methods employing dense prediction heads typically require high-resolution Bird’s Eye View (BEV) feature maps for conducting dense predictions on them. Considering computational complexity, some recent methods aim to establish global sparse and local dense prediction relationships [124, 131]. Meanwhile, certain detection methods focus on recovering 3D information from the detection process, for instance, the pioneering two-stage detection work Voxel-RCNN [113], which aggregates early features around voxels near instances using the Voxel ROI Pooling module to recover lost 3D information. Subsequent works have designed approaches based on the Voxel-RCNN paradigm, such as using corner point [180, 102, 104] or keypoints from [125, 105].
Numerous methods resort to auxiliary tasks to enhance the spatial features and provide implicit guidance for accurate 3D object detection, including IOU prediction to rectify the object confidence scores [151, 104, 103], object shape completion to complete object shapes from sparse point clouds [420, 162]. and object part estimation to gain 3D structure information by identifying the part information inside objects [372, 201].
@@
: In recent years, transformer [352, 353] has developed rapidly in Computer Vision and has shown amazing performance on numerous tasks. Therefore, many endeavors have been made to adapt Transformers to 3D object detection. Particularly, recent studies [253, 254] have confirmed the excellent robustness of transformer-based models, which will further advance research in the domain of safety perception for autonomous driving.
Compared with CNN, the query-key-value design and the self-attention mechanism make transformer modeling global relationships, resulting in a larger receptive field. However, the primary limitation for efficient application of Transformer-based models is the quadratic time and space complexity of the global attention mechanism. Hence, it’s critical to design specialized attention mechanisms for Transformer-based 3D object detectors. Transformer [351], DETR [355], and ViT [353] are the works that have most significantly influenced 3D transformer-based methods [127, 123, 126, 135, 181, 230, 223, 128]. They have each inspired subsequent 3D detection works in various aspects: the design of attention mechanisms, the architecture of encoders and decoders, and the development of patch-based inputs and architectures similar to visual transformers.
Inspired by transformer [351], VoTr [127] is the first work to incorporate transformer into a voxel-based backbone network, which is composed of sparse attention and sparse submanifold attention modules. Subsequent work [128] have continued to build on the foundation of voxel-transformer, further optimizing the temporal complexity of the attention mechanism. DETR [355] has inspired a range of networks to adopt an encoder-decoder structure akin to DETR’s. TransFusion [230] as a notable work, generates object queries from initial detections, applying cross-attention to LiDAR and image features within the Transformer decoder for 3D object detection. Meanwhile, many papers [181, 135, 123] try to explore and refine patch-based inputs mechanism from ViT [353] and the window attention mechanism from Swin Transformer [352], e.g. SST [123] and SWFormer [135] group local regions of voxels into patches, apply sparse regional attention, and then apply region shift to change the grouping. It is noteworthy that SEFormer [181] is the first to introduce object structure encoding into the transformer module.
@@ IV-C Point-based 3D object detection
Benefiting from the prosperity of point cloud in deep learning [421, 414, 422, 423], the Point-based 3D object detection inherits many framework and proposes to directly detect object from the raw points without preprocessing. Compared to voxel-based methods, the raw points maximally retains the original information, which is beneficial for fine-grained feature acquisition. Meanwhile, a series of point-based backbone works [414, 421] naturally provide a strong baseline for point-based methods. However, as of now, the performance of point-based methods is still influenced by two factors: the number of contextual points and the context radius used in feature learning, e.g. increasing the number of contextual points can provide fine-grain 3D information but significantly increases the model’s inference time. Similarly, reducing the context radius can achieve the same effect. Therefore, selecting appropriate values for these two factors enables the model achieve a balance between accuracy and speed. Specifically, to address the aforementioned issues, existing methods mostly focus on optimizing the two basic components of point-based 3D object detectors: 1) Point Cloud Sampling. 2) Feature Learning.
@@ IV-C1 Point Cloud Sampling
FPS (Farthest Point Sampling), originating from the work on PointNet++ [421], is a point cloud sampling method extensively utilized in point-based methods. It aims to select a set of representative points from the raw points, such that their mutual distances are maximized, thereby optimally covering the entire spatial distribution of the point cloud.
PointRCNN [111], a pioneering two-stage detector in point-based methods, utilize the PointNet++ [421] with multi-scale grouping as backbone network. In stage-1, it generates 3D proposals from point clouds in a bottom-up manner. The stage-2 network refines the proposals by combining semantic features and local spatial features. A similar framework proposed by [144]. Their model eliminates part of the background information with 2D image semantic segmentation.
However, existing methods relying on Farthest Point Sampling (FPS) still face several issues: 1) points irrelevant to detection also participate in the sampling process, leading to additional computational burden. 2) The distribution of points across different parts of an object is uneven, resulting in suboptimal sampling strategies. In subsequent works, design paradigms similar to FPS have been employed, with addressing the aforementioned issues, such as segmentation-guided background point filtering [186, 144], random sampling [424], feature space sampling [105], voxel-based sampling [107, 110], coordinate refinement [138] and ray-based grouping sampling [192].
IV-C2 Model Structure
The feature learning stage in point-based methods aims to extract discriminative feature representations from raw points. The neural network used in the feature learning phase should possess the following characteristics: 1) Invariance, where the point cloud backbone network should be insensitive to the ordering of input point clouds, 2) Local awareness, enabling the backbone network to perceive and model local regions and extract local features, 3) The ability to integrate contextual information, allowing the backbone network to extract features from both global and local contextual information.
Based on the aforementioned characteristics, a multitude of detectors have been designed for processing raw points. However, most methods can be categorized according to the core operators they utilize: 1) PointNet-based methods [111, 251, 195, 186, 425]. 2) Graph Neural Network-based methods [110, 426, 424, 427, 134]. 3) Transformer-based methods [138, 428].
@@ IV-C2 Model Structure
PointNet-based methods [111, 251, 195, 186, 425] primarily rely on the Set Abstraction [414] to perform downsampling on raw points, aggregation of local information, and integration of contextual information, while preserving the symmetry invariance of raw points. Point-RCNN [111], as the first two-stage work in point-based methods, achieved amazing performance at its time, yet it still faces the issue of high computational cost. Subsequent work [144, 186] has addressed this issue by introducing an additional semantic segmentation task during the detection process to filter out background points that contribute minimally to detection. Furthermore, some methods efforts have focused on resolving the issue of the uncontrolled receptive field in PointNet&PointNet++, such as through the use of GNN [161] or Transformer [138].
@@
GNNs (Graph Neural Networks) possess key elements such as an adaptive structure, dynamic neighborhood, the capability to construct both local and global contextual relationships, and robustness against irregular sampling. These characteristics naturally endow GNNs with an advantage in handling irregular point clouds. Point-GNN [110], a pioneering work, designs a one-stage graph neural network to predict object with an auto-registration mechanism, merging and scoring operation, which demonstrate the potential of using the graph neural network as a new approach for 3D object detection. Most graph-based point-based methods [110, 426, 424, 427, 429] aim to fully utilize contextual information. This motivation leaves room for further improvements in subsequent works [427, 429].
@@
Up to this point, a series of methods [126, 430, 428, 217, 138, 431] have explored the use of transformers for feature learning in point clouds and have achieved excellent results. Pointformer [138] introduced local and global attention modules for processing 3D point clouds. The local transformer module models interactions among points within local areas, aiming to learn contextually relevant regional features at the object level. The global transformer, on the other hand, focuses on learning context-aware representations at the scene level. Subsequently, the local-global Transformer combines local features with high-resolution global features to further capture dependencies between multi-scale representations. Group-free [428] directly utilize all points in the point cloud to compute features for each object candidate, where the contribution of each point is determined by an automatically learned attention module. Specifically, the authors adapted the Transformer to suit 3D object detection, enabling it to model both object-to-object and object-to-pixel relationships and extract object features without manual grouping. Moreover, by iteratively refining the spatial encoding of objects at different stages, the detection performance is further enhanced.
Point-based transformer directly process unstructured and unordered raw point clouds. This results in significantly higher computational complexity compared to structured voxel data. Consequently, the application of transformer-based methods in point-based methods is far less prevalent than in voxel-based methods.
@@ IV-D Point-Voxel based 3D object detection
Point-based methods offer high resolution and preserve the spatial structure of the original data, but they suffer from high computational complexity and inefficiency in handling sparse data. In contrast, voxel-based methods provide a structured data representation, enhancing computational efficiency and facilitating the application of conventional convolutional neural network techniques. However, due to the discretization process, they often lose fine spatial details. Driven by these issues, PV-based methods were developed. Point-voxel methods aim to leverage the fine-grained information capture capabilities of point-based methods and the computational efficiency of voxel-based methods. By integrating these methods, point-voxel based methods enable a more detailed processing of point cloud data, capturing both the global structure and micro-geometric details. This is critically important for safety perception in autonomous driving, as the accuracy of decisions made by autonomous driving systems depends on high-precision detection results.
The key goal of point-voxel methods is to enable feature interplay between voxels and points via point-to-voxel or voxel-to-point transformations. The idea that leverages point-voxel feature fusion in backbones has been explored by many works [116, 199, 146, 164, 166, 117, 165, 371, 432, 120, 433, 168, 139]. These methods fall into two categories: 1) Early Fusion. The early fusion methods fuses [116, 199, 146, 164, 166, 165] voxel features and point features within the backbone network. 2) Late Fusion. while the late fusion methods [117, 371, 432, 120, 433, 168, 139], typically a two-stage detection approach, uses voxel-based methods for initial proposal box generation, followed by sampling and refining key point features from the point cloud to enhance 3D proposals.
@@
Some methods [116, 199, 146, 163, 164, 165, 166] have explored using new convolutional operators to fuse voxel and point features, with PVCNN [116] potentially being the first work in this direction. In this approach, the voxel-based branch initially converts points into a low-resolution voxel grid and aggregates neighboring voxel features through convolution. Then, through a process called devoxelization, voxel-level features are transformed back to point-level features and fused with features obtained from the point-based branch. The point-based branch extracts features for each individual point. Since it does not aggregate neighboring information, the method can operate at a higher speed. Following closely, SPVCNN [199] which builds upon PVCNN, extends PVCNN to the domain of object detection. Other methods attempt to make improvements in different perspectives, such as auxiliary tasks [146, 163] or multiscale feature fusion [164, 165, 166].
@@
The methods in this series predominantly adopt a two-stage detection framework. Initially, voxel-based methods are employed to generate preliminary object proposals. This is followed by a refinement phase, where point-level features are leveraged for the precise delineation of detection boxes. Shi et al. proposed PV-RCNN [371], a milestone in PV-based methods. It utilizes SECOND [109] as the first-stage detector and proposes a second-stage refinement stage with a RoI grid pool for the fusion of keypoint features. Later works in this domain predominantly follow the aforementioned paradigm, focusing on advancements in second-stage detection. Notable developments include, attention mechanisms [167, 168, 139], scale-aware pooling [433] and point density-aware refinement modules [120].
PV-based methods simultaneously possess the computational efficiency of voxel-based approaches and the capability of point-based methods to capture fine-grained information. However, the construction of point-to-voxel or voxel-to-point relationships, along with the feature fusion of voxels and points, incurs additional computational overhead. Consequently, compared to voxel-based methods, PV-based methods can achieve better detection accuracy, but at the cost of increased inference time.
@@ IV-E Analysis: Accuracy, Latency, Robustness
In the autonomous driving sector, the development of LiDAR-based 3D object detection solutions is advancing rapidly. A series of works [1, 344, 345, 346, 347, 242, 2, 348, 349, 350, 269, 252, 270, 271, 272, 273, 274, 275] have comprehensively summarized the current technological roadmaps, such as the extensive review of LiDAR-based solutions by the Shanghai AI Lab and SenseTime Research [269]. However, there remains a lack of summarization and guidance from the perspective of safety perception and cost impact in autonomous driving. Therefore, in this section, following an analysis of the technological roadmaps and current state of LiDAR-based solutions, we intend to base our discussion on the fundamental principles of ‘Accuracy, Latency, and Robustness.’ This is aimed at providing guidance for the practical implementation of economically efficient and safe sensing in autonomous driving.

### v2

@@ IV-B2 Model Structure
There are three major types of neural networks in voxel-based methods: 1) 2D CNNs for processing BEV feature maps and pillars. 2) 3D Sparse CNNs for processing voxels. 3) Transformers for handling both voxels and pillars.
2D CNNs
2D CNNs are primarily used to detect 3D objects from a bird’s-eye view perspective, including processing BEV (Bird’s Eye View) feature maps and pillars [51, 93, 52, 91, 64]. Specifically, the 2D CNNs used for processing BEV feature maps often come from well-developed 2D object detection networks, such as Darknet [406], ResNet [407], FPN [401], and RPN [400]. One significant advantage of 2D CNNs compared to 3D CNNs is their faster speed. However, due to their difficulty in capturing spatial relationships and shape information, 2D CNNs typically exhibit lower accuracy.
3D Sparse CNNs
3D Sparse CNNs consist of two core operators: sparse convolution and submanifold convolution [408], which ensure that the convolutional operation is performed only on non-empty voxels. SECOND [43] implements efficient computation of sparse convolution [408] and submanifold convolution [409] operators to gain fast inference speed by constructing a hash table. It is followed by [46, 57, 47, 122, 39]. However, the limited receptive field of 3D Sparse CNNs, which leads to information truncation, restricts the model’s feature extraction capabilities. Meanwhile, the sparse representation of features makes it challenging for the model to capture fine-grained object boundaries and detailed information. To optimize these issues, main optimization strategies have emerged: 1) Expanding the model’s receptive field. Some methods [61, 60] extend the concept of large kernel convolution from 2D to 3D space or introduce additional downsampling layers in the model [46]. 2) Combining sparse and dense representations. Methods in this category typically utilize dense prediction heads to prevent information loss [57, 47, 42, 43, 302] or retrieve lost 3D information from the detection process [47, 86, 302, 37, 57], or they add additional auxiliary tasks to the model [302, 305, 38, 39, 77]. Methods employing dense prediction heads typically require high-resolution Bird’s Eye View (BEV) feature maps for conducting dense predictions on them. Considering computational complexity, some recent methods aim to establish global sparse and local dense prediction relationships [56].
@@
Due to the amazing performance of transformers [371, 372], many efforts have been made to adapt Transformers to 3D object detection. Particularly, recent studies [124, 125] have confirmed the excellent robustness of transformer-based models, which will further advance research in the domain of safety perception for autonomous driving. Compared with CNNs, the query-key-value design and the self-attention mechanism allow transformers to model global relationships, resulting in a larger receptive field. However, the primary limitation for efficiently applying Transformer-based models is the quadratic time and space complexity of the global attention mechanism. Hence, designing specialized attention mechanisms for Transformer-based 3D object detectors is critical. Transformer [370], DETR [374], and ViT [372] are the works that have most significantly influenced 3D transformer-based methods [58, 55, 64, 91, 108, 101, 59]. They have each inspired subsequent 3D detection works in various aspects: the design of attention mechanisms, the architecture of encoders and decoders, and the development of patch-based inputs and architectures similar to visual transformers. Inspired by transformer [370], VoTr [58] is the first work to incorporate a transformer into a voxel-based backbone network, composed of sparse attention and sparse submanifold attention modules. Subsequent works [59] have continued to build on the foundation of voxel-transformer, further optimizing the temporal complexity of the attention mechanism. DETR [374] has inspired a range of networks to adopt an encoder-decoder structure akin to DETR’s. TransFusion [108] is a notable work that generates object queries from initial detections, applying cross-attention to LiDAR and image features within the Transformer decoder for 3D object detection. Meanwhile, many papers [91, 64, 55] are exploring and refining the patch-based input mechanism from ViT [372] and the window attention mechanism from Swin Transformer [371], e.g., SST [55] and SWFormer [64] group local regions of voxels into patches, apply sparse regional attention, and then apply region shift to change the grouping. Notably, SEFormer [91] is the first to introduce object structure encoding into the transformer module.
@@ IV-C Point-based 3D object detection
Unlike voxel-based methods, point-based methods retain the original information to the maximum extent, facilitating fine-grained feature acquisition. However, the performance of point-based methods is still affected by two crucial factors: 1) the number of contextual points in the point cloud sampling stage and 2) the context radius used in the point-based backbone. These factors significantly impact the speed and accuracy of point-based methods, including the detection of small objects, which is critical for safety considerations. Therefore, optimizing these two factors is paramount, based on existing literature. In this regard, we primarily focus on elucidating 1) Point Cloud Sampling and 2) Point-based Backbone.
@@ IV-C1 Point Cloud Sampling
As an extensively utilized method, FPS (Farthest Point Sampling) aims to select a set of representative points from the raw points, such that their mutual distances are maximized, thereby optimally covering the entire spatial distribution of the point cloud.
PointRCNN [45], a pioneering two-stage detector in point-based methods, utilizes the PointNet++ [410] with multi-scale grouping as the backbone network. In the first stage, it generates 3D proposals from point clouds in a bottom-up manner. The second stage network refines the proposals by combining semantic features and local spatial features. However, existing methods relying on FPS still face several issues: 1) Points irrelevant to detection also participate in the sampling process, leading to additional computational burden. 2) The distribution of points across different parts of an object is uneven, resulting in suboptimal sampling strategies. Subsequent works have attempted various optimization strategies, such as segmentation-guided background point filtering [94], random sampling [295], feature space sampling [40], voxel-based sampling [113, 44], coordinate refinement [66], and ray-based grouping sampling [95].
IV-C2 Point-based Backbone
The feature learning stage in point-based methods aims to extract discriminative feature representations from raw points. The neural network used in the feature learning phase should possess the ability of to local be awareness locally aware and integrating to integrate contextual information. Based on the aforementioned motivations, a multitude of detectors have been designed for processing raw points. However, most methods can be categorized according to the core operators they utilize: 1) PointNet-based methods [45, 98, 94, 297]. 2) Graph Neural Network-based methods [44, 294, 295, 411, 63]. 3) Transformer-based methods [66, 412].
@@ IV-C2 Point-based Backbone
PointNet-based methods [45, 98, 94, 297] primarily rely on the Set Abstraction [405] to perform downsampling on raw points, aggregation of local information, and integration of contextual information, while preserving the symmetry invariance of the raw points. Point-RCNN [45], as the first two-stage work in point-based methods, achieved amazing performance at its time; however, it still faces the issue of high computational cost. Subsequent work [70, 94] has addressed this issue by introducing an additional semantic segmentation task during the detection process to filter out background points that contribute minimally to detection. Furthermore, some efforts have focused on resolving the issue of the uncontrolled receptive field in PointNet & PointNet++, such as through the use of GNN [80] or Transformer [66] techniques.
@@
GNNs (Graph Neural Networks) possess key elements such as an adaptive structure, dynamic neighborhood, the capability to construct both local and global contextual relationships, and robustness against irregular sampling. These characteristics naturally endow GNNs with an advantage in handling irregular point clouds. Point-GNN [44], a pioneering work, designs a one-stage graph neural network to predict objects with an auto-registration mechanism, merging, and scoring operations, which demonstrate the potential of using graph neural networks as a new approach for 3D object detection. Most graph-based, point-based methods [44, 294, 295, 411, 298] aim to fully utilize contextual information. This motivation has led to further improvements in subsequent works [411, 298].
@@
Up to this point, a series of methods [413, 412, 66, 414] have explored the use of transformers for feature learning in point clouds, achieving excellent results. Pointformer [66] introduced local and global attention modules for processing 3D point clouds. The local transformer module models interactions among points within local areas, with the aim of learning contextually relevant regional features at the object level. The global transformer, on the other hand, focuses on learning context-aware representations at the scene level. Subsequently, the local-global Transformer combines local features with high-resolution global features to further capture dependencies between multi-scale representations. Group-free [412] adapted the Transformer to suit 3D object detection, enabling it to model both object-to-object and object-to-pixel relationships and to extract object features without manual grouping. Moreover, by iteratively refining the spatial encoding of objects at different stages, the detection performance is further enhanced. Point-based transformers directly process unstructured and unordered raw point clouds, which results in significantly higher computational complexity compared to structured voxel data.
@@ IV-D Point-Voxel based 3D object detection
Point-voxel methods aim to leverage the fine-grained information capture capabilities of point-based methods and the computational efficiency of voxel-based methods. By integrating these methods, point-voxel based methods enable a more detailed processing of point cloud data, capturing both the global structure and micro-geometric details. This is critically important for safety perception in autonomous driving, as the accuracy of decisions made by autonomous driving systems depends on high-precision detection results.
The key goal of point-voxel methods is to enable feature interplay between voxels and points via point-to-voxel or voxel-to-point transformations. The idea of leveraging point-voxel feature fusion in backbones has been explored by many works [49, 97, 72, 82, 84, 50, 83, 308, 415, 53, 381, 86, 67]. These methods fall into two categories: 1) Early Fusion. Early fusion methods [49, 97, 72, 82, 84, 83] fuse voxel features and point features within the backbone network. 2) Late Fusion. Late fusion methods [50, 308, 415, 53, 381, 86, 67] typically employ a two-stage detection approach, using voxel-based methods for initial proposal box generation, followed by sampling and refining key point features from the point cloud to enhance 3D proposals.
@@
Some methods [49, 97, 72, 82, 83, 84] have explored using new convolutional operators to fuse voxel and point features, with PVCNN [49] potentially being the first work in this direction. In this method, the voxel-based branch initially converts points into a low-resolution voxel grid and aggregates neighboring voxel features through convolution. Then, the voxel-level features are transformed back into point-level features and fused with the features obtained from the point-based branch. Following closely, SPVCNN [97], which builds upon PVCNN, extends PVCNN to the domain of object detection. Other methods attempt to make improvements from other perspectives, such as auxiliary tasks [72] or multi-scale feature fusion [82, 83, 84].
@@
The methods in this series predominantly adopt a two-stage detection framework. Initially, voxel-based methods are employed to generate preliminary object proposals. This is followed by a refinement phase, where point-level features are leveraged for the precise delineation of detection boxes. As a milestone in PV-based methods, PV-RCNN [308] utilizes SECOND [43] as the first-stage detector and proposes a second-stage refinement stage with a RoI grid pool for the fusion of keypoint features. Subsequent works have followed the aforementioned paradigm, focusing on advancements in second-stage detection. Notable developments include the use of attention mechanisms [85, 86, 67], scale-aware pooling [381], and point density-aware refinement modules [53].
PV-based methods simultaneously possess the computational efficiency of voxel-based approaches and the capability of point-based methods to capture fine-grained information. However, constructing point-to-voxel or voxel-to-point relationships, along with the feature fusion of voxels and points, incurs additional computational overhead. Consequently, compared to voxel-based methods, PV-based methods can achieve better detection accuracy and robustness, but at the cost of increased inference time.
@@ IV-E Analysis: Accuracy, Latency, Robustness
In the autonomous driving sector, the development of LiDAR-only 3D object detection solutions is advancing rapidly. A series of works [4, 126, 123, 130, 131, 1] have comprehensively summarized the current technological roadmaps, such as the extensive review of LiDAR-only solutions by the Shanghai AI Lab and SenseTime Research [126]. However, there is a lack of summarization and guidance from the perspective of safety perception and cost impact in autonomous driving. Therefore, in this section, following an analysis of the technological roadmaps and the current state of LiDAR-only solutions, we intend to base our discussion on the fundamental principles of ‘Accuracy, Latency, and Robustness.’ It aims to guide the practical implementation of economically efficient and safe sensing in autonomous driving.